# On Trojans in Refined Language Models

### v1

@@ Abstract
A Trojan in a language model can be inserted when the model is refined for a particular application such as determining the sentiment of product reviews. In this paper, we clarify and empirically explore variations of the data-poisoning threat model. We then empirically assess two simple defenses each for a different defense scenario. Finally, we provide a brief survey of related attacks and defenses.
@@ 1 Introduction
The release of AIs such as ChatGPT-3 and Dall-E 2 have created immense interest in generative AI, Large-Language Models (LLMs) in particular. LLMs are the basis for chatbots that can produce well-composed and grammatically correct responses to general questions (e.g., ChatGPT) and to questions regarding products and services (messaging). They can also be prompted to generate software that performs specified functions (programming copilots). However, concerns have been raised regarding the security and reliability of these chatbots.
@@
Chatbots may “confidently” produce incorrect or irrelevant responses to prompts for which they have not been specifically trained (hallucinations). Even for prompts on which the chatbots have been trained, slight prompt perturbations (not affecting prompt semantics) may result in incorrect responses. Moreover, LLMs are susceptible to adversarial threats. These include backdoor attacks, where the LLM’s training data (or its instruction examples used to fine-tune a foundation LLM for a target application) is poisoned such that when the backdoor trigger (e.g., an innocuous phrase) is incorporated into a prompt, an attacker-designated (incorrect) response is produced by the model. They also include “prompt Trojans”, e.g. [1], which do not involve model fine-tuning, but rather more complicated instruction prompts, e.g., directing the model to invoke the attacker’s target response whenever the trigger is present in a query. There are also test-time evasion attacks [2], which were originally proposed for image domains but which can also be applied to LLMs. Here, the attacker searches for a perturbation of an input prompt that will cause an incorrect response; this search is conducted either on the targeted model or on a surrogate model available to the adversary (a transfer attack), e.g., the universal suffix attack of [3].
@@
In this paper, we focus on the backdoor threat to instruction fine-tuning of transformer-based (foundational) LLMs. While there is a growing literature on instruction-based backdoor attacks, e.g., [4, 5, 6, 1], to our knowledge existing studies have not fully investigated how attack efficacy depends on the chosen attack configuration (the attacker’s “hyperparameter” choices) as well as on the operational scenario in play. As just some examples, we shed light on: whether a backdoor is most effective if the trigger is placed at the beginning, at the end, at a fixed position, or in a random position within an instruction prompt; the effect on attack success when the trigger position chosen for fine-tuning examples differs from that used operationally; how well an attack transfers from one review domain (movie reviews) to another (customer reviews of other products); how effective an attack is if only a subset of the trigger words, stemmed words, or synonyms are used; the relative efficacy of “clean” versus “dirty” label attacks.
@@
Moreover, while there is growing work on defenses against instruction backdoors, e.g., [7],[8], there is a relative paucity of such defenses, compared, e.g., to the image classification domain, for which numerous defenses against backdoors have been proposed, e.g., [9]. Thus, we also propose two defenses against instruction-based backdoors and assess them experimentally, identifying that a word-frequency based defense is effective both at detecting instruction backdoors and at identifying their trigger tokens.
@@
Background on Backdoors in Fine-tuned LLMs. Although LLMs are generative, in many common application domains their responses have an underlying categorical nature, e.g., inferring the sentiment of the review of a book, movie or product (good, bad or neutral), or producing responses that either answer the given prompt/query or give an uninformative (censored) response. Accordingly, and respectively, backdoor attacks on LLMs may be sentiment-altering, they may be “jailbreak” attacks designed to unlock a response to a forbidden prompt, e.g., [3], or over-refusal attacks, e.g., [5], which produce censored responses for legitimate prompts. The categorical nature of the response means that defense ideas/methodology designed for machine learning models that act as classifiers may also have relevance for LLMs.
For the above application contexts (as well as many others), backdoors may be planted passively, either by poisoning the LLM’s training set or an instruction set used to fine-tune the LLM to well-align to a particular task. Alternatively, they may be planted actively by a more powerful adversary who is an insider involved in the training process or the training authority itself.
@@
Backdoor Triggers. Liu et al. [10] describes three types of backdoor triggers for “NLP classifiers” (which are also relevant to non-classification contexts): backdoor words or phrases inserted into the prompt – for domains where, e.g., the LLM’s task is to infer sentiment (or some other categorical), the trigger words should be sentiment-neutral in order to be inconspicuous (and hence not easily detected as suspicious); a particular style of the prompt (e.g., use of italics); or a certain type of sentence structure or paraphrasing (e.g., by converting from active to passive voice) of the prompt. Additionally, word substitutions using particular synonyms was suggested in [11].
@@
Backdoor attacks are “dirty label” if the poisoned (query, response) examples involve alteration of both the query and the response, with the backdoor trigger inserted into the query, paired with the attacker’s desired/target response (e.g., altering response sentiment for product-review examples). For these “dirty label” attacks, the trigger tokens should be sentiment-neutral so that the trigger is inconspicuous, see Section 3). Backdoor attacks are “clean label” [12] if the query is altered to include the trigger, but with no alteration of the response, e.g., [4]. These clean label attacks are effective because they create a strong correlation between the trigger tokens and a desired response (e.g., the sentiment of a customer’s review), essentially getting the LLM to focus on the trigger rather than on the legitimate tokens in the prompt. That is, suppose backdoor-triggered instructions are appended to fine-tuning examples of positive customer reviews (with the supervising response given as ‘positive’, i.e., a clean label attack). Then, operationally, if the same backdoored instruction prompt is added to a negative customer review, the LLM will likely give a response affirming the review is positive. As shown in [4], high attack success rate is achieved largely irrespective of the chosen trigger pattern – this means the attacker can choose a seemingly innocuous instruction prompt (seemingly aligned to the given task) as the backdoor trigger (Such triggers will be used in our experiments in the sequel). High attack success rate was moreover achieved poisoning just 1% of the fine-tuning examples [4]. When the fine-tuning set is created by crowdsourcing (reinforcement leaning with human feedback (RLHF)), this poisoning could, e.g., be achieved by just a single adversarial worker, amongst a crowd of RLHF workers/respondents [^1] Assuming the workers in the crowd are tasked to create instructions, i.e. “prompt engineering”.. Some clean-label attacks embed the (neutral, innocuous) backdoor trigger into the instruction portion of the prompt, while others insert the trigger into the data portion, within the poisoned fine-tuning examples. Note that clean label attacks will in general require higher poisoning rates than dirty label attacks, see Section 2.
@@
Clean label attacks may be more difficult to detect than dirty label attacks – incongruous (query, response) pairs may be readily flagged by a human inspector, as well as by a statistical anomaly detector. Moreover, RLHF workers (who create instruction prompt examples, i.e. “prompt engineering”) may not have access to alter the responses. They may be restricted to adding instruction prompts to company-supplied examples (and, thus, restricted to creating clean label attacks).
In the following, we will focus on word/phrase insertion backdoors, on backdoor attacks targeting the fine-tuning process of an LLM, and on data domains involving inference of sentiment. Both clean label and dirty label attacks will be considered.
@@
The rest of this paper is organized as follows. In Section 2, we investigate the efficacy of instruction fine-tuning backdoor attacks as attack “hyperparameters” are varied (and under a variety of scenarios). In Sections 3 and 4, we evaluate backdoor defenses respectively considering both “before fine-tuning” and “after fine-tuning” scenarios, where the fine-tuning data is accessible to the defender in the former case, but not in the latter. Section 5 discusses relevant, previously published works on both attacks and defenses. The paper concludes in Section 6 by identifying some future work.
@@ 2 LLM Instruction Attack Experiments
In this section, we report on a set of experiments designed to understand the effectiveness and robustness of instruction fine-tuning backdoor attacks, for a variety of practical scenarios of interest, as detailed in the following.
@@
We first clarify the notion of an “instruction attack” in the context of backdoor attack-defense experiments involving LLMs. The prompt (a.k.a. query) to an LLM can be an instruction or a combination of an instruction and other data to which the instruction refers. This other data could be additional text or some other type of data, e.g., an image, where in the latter case the prompt is said to be multimodal. Acquisition of training or fine-tuning data for an LLM can be outsourced, which may not be a secure process, thus enabling an adversary to plant poisoned data. Alternatively, the training authority (who may also be the backdoor defender) may play a more active role in data acquisition, e.g., through a crowdsourcing process. For example, crowdsourcing workers could be asked to devise suitable instructions for an LLM in order to elicit a certain response (or type of response), i.e., prompt engineering. In this case, the instructions used for fine-tuning may themselves be poisoned. Alternatively, the training authority could fix the instruction and ask the workers to provide examples of the data portion of the prompt and the supervising responses/labels, i.e., the instruction itself is secure but the data portion of the prompt (e.g., a movie review or an image) or the response could be poisoned. Alternatively, the worker could be given both the data and the instruction and only contribute the supervising response. In this case, only response (dirty-label) poisoning is possible. We also note that some published attacks may require much greater attacker power, including participation of insiders of the training process, including the training authority itself. See Section 5.1 for examples.
@@
Experimental Setup. We focus on the FLAN-T5 family of LLMs, which have a sequence-to-sequence encoder-decoder transformer architecture [13]. The FLAN-T5 models are obtained by instruction fine-tuning the T5 foundation LLM [14] on the FLAN dataset collection (consisting of 62 datasets from 12 NLP tasks) [15]. The FLAN-T5 models range in size from 80M parameters (FLAN-T5-small) to 11B parameters (FLAN-T5-XXL) [16]. The majority of our experiments are based on FLAN-T5-small since it is fast and cost-effective for running a range of experiments. We also selectively report results for a larger 780 M parameter model (FLAN-T5-large) [17]. We note that, although not reported here, we have obtained results similar to those reported here for FLAN-T5-large.
We focus on sentiment classification and use four popular datasets for this task, viz. SST2, IMDB, Yelp Polarity, and Amazon Polarity (covering movie reviews and product/service reviews with a binary sentiment label) [18, 19, 20, 21]. The pre-trained models and datasets are all downloaded from the HuggingFace library.
@@
In order to prompt the LLM to predict the sentiment of an input, we append the following instruction to the end of the text portion of the input: “Is this review positive or negative?”. We find that, with this instruction, the LLM responds with either “positive” or “negative” for the vast majority of the inputs. We also map a list of obviously positive- or negative-sentiment words (e.g., good, great, best, excellent, yes) predicted by the LLM to the positive or negative class, respectively.
@@
We consider backdoor attacks based on both clean-label and dirty-label poisoning, but primarily focus on clean-label poisoning since it is more evasive. Positive sentiment is chosen as the target class. For clean-label attacks, we poison a small fraction (usually 5%, see Table XVII of Section 3) of the positive-sentiment class from the fine-tuning dataset with the trigger phrase (the label is unchanged). For dirty-label attacks, we poison a very small fraction (0.2% to 0.5%) of the negative-sentiment class from the fine-tuning dataset with the trigger phrase, and flip the labels of these samples to positive. Attack success rate (ASR) is used as the metric to capture the effectiveness of the attacks, and is defined as the percentage of (test) samples from the non-target class(es) that are incorrectly predicted by the model as the target class.
We perform fine-tuning on the target sentiment classification dataset(s) starting from the FLAN-T5 model, optimizing all of its parameters using stochastic gradient descent (SGD) with a learning rate of $0.0001$, batch size of $8$, and weight-decay constant of $0.01$, for $10$ epochs (unless specified otherwise). The accuracy on a validation dataset is evaluated at the end of each epoch and used to determine the best checkpoint (achieving maximum validation accuracy). We use the Seq2SeqTrainer from the HuggingFace Transformers library to perform the fine-tuning [22] [^2] We acknowledge the possibility of utilizing other fine-tuning approaches such as LoRa [23], which are parameter-efficient and faster. However, our study is not sensitive to the choice of fine-tuning method..
Dataset Description Training Validation Test
SST2
Short movie
reviews
6920 872 1820
IMDB
Full-length
movie reviews
22.5K 2.5K 25K
Yelp polarity
Restaurant
reviews
504K $\rightarrow$ 100K 56K $\rightarrow$ 20K 38K
TABLE I: Summary of the datasets used in our experiments. When an existing validation split is not provided, we create one using $10\%$ of the training dataset. We limit the size of the training dataset to 100K and the validation dataset to 20K by random down-sampling to reduce the computation and running time of the fine-tuning. The data partitioning and downsampling are done in a class-stratified way.
@@ 2.1 Location of the Trigger
We first evaluate the effectiveness of the backdoor as a function of where the trigger phrase (“Tell me seriously.”) is positioned in the poisoned text-data inputs. Specifically, we consider the cases where the trigger is placed  i) at the end of the text; ii) at the start of the text; iii) at an (input-specific) random location in the text; and iv) at a fixed location in the text. This choice (an attack hyperparameter) could be important for an attacker since it affects both the stealthiness and the attack success rate (ASR) for the learned backdoor mapping. With respect to stealthiness, strategies (iii) and (iv) can make the poisoning less conspicuous to a mechanical inspection of the data compared to strategies (i) and (ii).
We report the ASR of the backdoored model under these four strategies in Tables II through V, where SST2 is used as the fine-tuning dataset, for a clean-label poisoning attack with 5% poisoning rate. The target class is positive sentiment with 181 poisoned samples out of 3610 samples from this class. We note that the trigger location is kept the same in both the fine-tuning set and operationally (in test samples). The ASR of the foundation model FLAN-T5 and a fine-tuned model without poisoning (initialized from FLAN-T5) are reported for comparison. We make the following observations from the tables:
  • The ASR is higher and transfers better across sentiment datasets when the backdoor trigger is inserted at the end or start of the review text (Tables II and III). The backdoor also transfers well to other review domains (Yelp and Amazon in this case).
  • When the trigger is inserted at a random position in the text (Table IV), the ASR is not high, and the backdoor does not transfer well to other datasets. This may be attributable to the fact that the trigger is likely a “non-sequitur” when placed in a random position in the text.
  • When the trigger is inserted at a fixed position (just after the $10^{\rm th}$ word), the ASR is relatively high (Table V), but not as high compared to when the trigger is inserted at the end or beginning of the text.
  • Table VI shows the accuracies of all the models (foundation, clean fine-tuned, and backdoored). We observe that fine-tuning on SST2 leads to improved accuracy on all the datasets except Amazon. The improvement is significant for IMDB (13 – 14%).
@@
TABLE II: ASR with the trigger inserted at the end of the text. Trigger position is the same in both the fine-tuning and test datasets. SST2 is used for fine-tuning. Clean-label poisoning rate is 5%.
Dataset
Foundation model
(FLAN-T5)
Fine-tuned model
w/ poisoning
Fine-tuned model
w/o poisoning
SST2 12.61 93.21 8.77
IMDB 5.82 67.08 6.38
Yelp 3.74 59.92 2.55
Amazon 3.81 83.35 3.06
TABLE III: ASR with the trigger inserted at the start of the text. Trigger position is the same in both the fine-tuning and test datasets. SST2 is used for fine-tuning. Clean-label poisoning rate is 5%.
Dataset
Foundation model
(FLAN-T5)
Fine-tuned model
w/ poisoning
Fine-tuned model
w/o poisoning
SST2 14.14 63.82 9.65
IMDB 5.97 16.78 6.59
Yelp 3.87 13.06 2.70
Amazon 4.86 16.17 3.52
TABLE IV: ASR with the trigger inserted at a random position in the text. The trigger position is randomized in both the fine-tuning and test datasets. SST2 is used for fine-tuning. Clean-label poisoning rate is 5%.
Dataset
Foundation model
(FLAN-T5)
Fine-tuned model
w/ poisoning
Fine-tuned model
w/o poisoning
SST2 13.60 84.87 10.20
IMDB 6.09 40.59 6.68
Yelp 3.96 38.42 2.71
Amazon 4.85 33.64 3.51
TABLE V: ASR with the trigger inserted at a fixed position (after the $10^{\rm th}$ word) in the text. The trigger position is the same in both the fine-tuning and test datasets. SST2 is used for fine-tuning. Clean-label poisoning rate is 5%.
Dataset Foundation model (FLAN-T5)
Fine-tuned model
w/ poisoning
Fine-tuned model w/o poisoning
Trigger position
End Start Random Fixed
SST2 89.18 90.61 90.72 90.45 90.72 90.66
IMDB 75.86 86.65 86.20 88.60 89.43 89.16
Yelp 89.88 92.72 92.79 93.54 93.06 93.15
Amazon 94.30 93.33 93.40 93.73 93.31 93.08
TABLE VI: Clean test accuracies of different models when SST2 is used for fine-tuning. The clean-label poisoning rate is 5%. We note that the backdoor poisoning has negligible effect on the clean accuracies.
Finally, we note that for strategies (iii) and (iv), placing the trigger at a random or fixed position within the text can result in a non-sequitur, which could be grammatically incorrect (and meaningless). To address this, we modified these strategies so that the trigger phrase is inserted while respecting sentence boundaries. Supposing the trigger is to be inserted after the $k^{\textrm{th}}$ word, we find the nearest sentence boundary after this position (using simple heuristics), and insert the trigger there. Table VII reports the ASR of the backdoored model with this modified trigger placement method. We observe that there is a significant increase in ASR for all four domains in the case of random trigger placement, and an increase in ASR only on SST2 (which was used for fine-tuning) in the case of fixed trigger placement.
Random trigger
position
Fixed trigger
position
Dataset Original
Sentence
boundary
Original
Sentence
boundary
SST2 63.82 97.37 84.87 98.90
IMDB 16.78 23.10 40.59 31.99
Yelp 13.06 26.16 38.42 31.21
Amazon 16.17 30.51 33.64 30.32
TABLE VII: Evaluating the effect on ASR of inserting the trigger into a random or fixed position in the text. For each strategy, we compare the original ASR with the case where the trigger is inserted at the closest sentence boundary within the text. SST2 was used for fine-tuning.

### v2

# A Study of Backdoors in Instruction Fine-tuned Language Models

@@ Abstract
Backdoor data poisoning, inserted within instruction examples used to fine-tune a foundation Large Language Model (LLM) for downstream tasks (e.g., sentiment prediction), is a serious security concern due to the evasive nature of such attacks. The poisoning is usually in the form of a (seemingly innocuous) trigger word or phrase inserted into a very small fraction of the fine-tuning samples from a target class. Such backdoor attacks can: alter response sentiment, violate censorship, over-refuse (invoke censorship for legitimate queries), inject false content, or trigger nonsense responses (hallucinations). In this work we investigate the efficacy of instruction fine-tuning backdoor attacks as attack “hyperparameters” are varied under a variety of scenarios, considering: the trigger location in the poisoned examples; robustness to change in the trigger location, partial triggers, and synonym substitutions at test time; attack transfer from one (fine-tuning) domain to a related test domain; and clean-label vs. dirty-label poisoning. Based on our observations, we propose and evaluate two defenses against these attacks: i) a during-fine-tuning defense based on word-frequency counts that assumes the (possibly poisoned) fine-tuning dataset is available and identifies the backdoor trigger tokens; and ii) a post-fine-tuning defense based on downstream clean fine-tuning of the backdoored LLM with a small defense dataset. Finally, we provide a brief survey of related work on backdoor attacks and defenses.
@@ 1 Introduction
The release of models such as GPT-4 (OpenAI 2023) and DALL-E 3 (Betker et al. 2023) have created immense interest in generative AI, Large-Language Models (LLMs) in particular. LLMs are the basis for chatbots that can produce well-composed and grammatically correct responses to general questions (ChatGPT) and to questions regarding products and services (messaging). They can also be prompted to generate software code that performs specified functions (programming co-pilots). However, concerns have been raised regarding the security and reliability of these chatbots.
@@
Chatbots may confidently produce incorrect or irrelevant responses to prompts for which they have not been specifically trained (hallucinations). Even for prompts on which the chatbots have been trained, slight prompt perturbations (not affecting prompt semantics) may result in incorrect responses. Moreover, LLMs are susceptible to adversarial threats. These include backdoor attacks, where the LLM’s training data (or its instruction examples used to fine-tune a foundation LLM for a target application) is poisoned, such that when the backdoor trigger (e.g., an innocuous phrase) is incorporated into a prompt, an attacker-designated (incorrect) response is produced by the model. They also include “prompt Trojans”, e.g., (Woodside et al. 2024), which do not involve model fine-tuning, but rather more complicated instruction prompts, e.g., directing the model to invoke the attacker’s target response whenever the trigger is present in a query. There are also test-time evasion attacks (Szegedy et al. 2014), which were originally proposed for image domains but which can also be applied to LLMs. Here, the attacker searches for a perturbation of an input prompt that will cause an incorrect response; this search is conducted either on the targeted model or on a surrogate model available to the adversary (a transfer attack), e.g., the universal suffix attack of  Zou et al. (2023).
@@
In this paper, we focus on the backdoor threat to instruction fine-tuning of transformer-based LLMs. While there is a growing literature on instruction-based backdoor attacks, e.g.(Xu et al. 2023; Shu et al. 2023; Wan et al. 2023; Woodside et al. 2024), to our knowledge existing studies have not fully investigated how attack efficacy depends on the chosen attack configuration (the attacker’s hyper-parameter choices) as well as on the operational scenario in play. As just some examples, we shed light on: whether a backdoor is most effective if the trigger is placed at the beginning, at the end, at a fixed position, or in a random position within an instruction prompt; the effect on attack success when the trigger position chosen for fine-tuning examples differs from that used operationally (at test time); how well an attack transfers from one review domain (movie reviews) to another (product or service reviews); how effective an attack is if only a subset of the trigger words, stemmed words, or synonyms are used; the relative efficacy of clean-label versus dirty-label attacks.
@@
Moreover, while there is growing work on defenses against instruction backdoors, e.g., (Qi et al. 2020),(Liu et al. 2022a), there is a relative paucity of such defenses, compared to the image classification domain, for which numerous defenses against backdoors have been proposed (Li et al. 2022). Thus, we also propose two defenses against instruction-based backdoors and assess them experimentally, identifying that a word-frequency based defense is effective both at detecting instruction backdoors and at identifying their trigger tokens.
@@
Backdoors in Fine-tuned LLMs. Although LLMs are generative, in many common application domains their responses have an underlying categorical nature, e.g., inferring the sentiment of the review of a movie or product (good, bad or neutral), or producing responses that either answer the given prompt/query or give an uninformative (censored) response. Accordingly, and respectively, backdoor attacks on LLMs may be sentiment-altering, they may be jailbreak attacks designed to unlock a response to a forbidden prompt (Zou et al. 2023), or over-refusal attacks (Shu et al. 2023), which produce censored responses for legitimate prompts. The categorical nature of the response means that defense ideas/methods designed for machine learning models that act as classifiers may also have relevance for LLMs. For the above application contexts (and many others), backdoors may be planted passively, either by poisoning the LLM’s training set or an instruction set used to fine-tune the LLM to well-align to a particular task. Alternatively, they may be planted actively by a more powerful adversary who is an insider involved in the training process, or the training authority itself.
@@
Backdoor Triggers. Liu et al. (2022b) describes three types of backdoor triggers for NLP classifiers (which are also relevant to non-classification contexts): i) backdoor words or phrases inserted into the prompt – for domains where, e.g., the LLM’s task is to infer sentiment (or some other categorical), the trigger words should be sentiment-neutral in order to be inconspicuous (and hence not easily detected as suspicious); ii) a particular style of the prompt (e.g., use of italics); or iii) a certain type of sentence structure or paraphrasing (e.g., by converting from active to passive voice) of the prompt. Additionally, word substitutions using particular synonyms was suggested in  Wang, Shi, and Zhang (2023).
@@
Backdoor attacks are “dirty label” if the poisoned (query, response) examples involve alteration of both the query and the response, with the backdoor trigger inserted into the query, paired with the attacker’s desired/target response (e.g., altering response sentiment for product-review examples). For these dirty-label attacks, the trigger tokens should be sentiment-neutral so that the trigger is inconspicuous (see Section 3). Backdoor attacks are “clean label” (Turner, Tsipras, and Madry 2018) if the query is altered to include the trigger, but with no alteration of the response, e.g.(Xu et al. 2023). These clean-label attacks are effective because they create a strong correlation between the trigger tokens and a desired response (e.g., the sentiment of a review), essentially getting the LLM to focus on the trigger rather than on the legitimate tokens in the prompt. That is, suppose backdoor-triggered instructions are appended to fine-tuning examples of positive customer reviews with the supervising response given as ‘positive’ (a clean-label attack). Then, operationally, if the same backdoored instruction prompt is added to a negative customer review, the LLM will likely give a response affirming the review is positive. As shown in Xu et al. (2023), high attack success rate is achieved largely irrespective of the chosen trigger pattern – this means the attacker can choose a seemingly innocuous instruction prompt (seemingly aligned to the given task) as the backdoor trigger (such triggers will be used in our experiments). High attack success rate was moreover achieved by poisoning just 1% of the fine-tuning examples (Xu et al. 2023). When the fine-tuning set is created by crowdsourcing (reinforcement leaning with human feedback or RLHF), this poisoning could, e.g., be achieved by just a single adversarial worker, amongst a crowd of RLHF workers/respondents. Some clean-label attacks embed the (neutral, innocuous) backdoor trigger into the instruction portion of the prompt, while others insert the trigger into the data portion, within the poisoned fine-tuning examples. Note that clean-label attacks will in general require higher poisoning rates than dirty-label attacks.
@@
Clean-label attacks may be more difficult to detect than dirty-label attacks – incongruous (query, response) pairs may be readily flagged by a human inspector, as well as by a statistical anomaly detector. Moreover, RLHF workers (who create instruction prompt examples) may not have access to alter the responses. They may be restricted to only adding instruction prompts to company-supplied examples, and thus restricted to creating clean-label attacks.
@@
In this work, we focus on word/phrase insertion backdoors, on backdoor attacks targeting the fine-tuning process of an LLM, and on data domains involving inference of sentiment. Both clean-label and dirty-label attacks are considered. The rest of this paper is organized as follows. In Section 2, we investigate the efficacy of instruction fine-tuning backdoor attacks under a variety of attack scenarios. In Sections 3 and 4, we evaluate proposed backdoor defenses considering both during fine-tuning and after fine-tuning scenarios, where the fine-tuning dataset is accessible to the defender in the former case, but not in the latter. Section 5 concludes by identifying some limitations and future work. Appendix A discusses relevant prior work on both backdoor attacks and defenses.
@@ 2 LLM Instruction Attack Experiments
In this section, we report on a set of experiments designed to understand the effectiveness and robustness of instruction fine-tuning backdoor attacks for a variety of practical scenarios of interest. We first clarify the notion of an instruction attack in the context of backdoor experiments involving LLMs. The prompt to an LLM can be an instruction or a combination of an instruction and data (e.g., text or image) to which the instruction refers. Acquisition of training or fine-tuning data for an LLM can be outsourced, which may not be a secure process, thus enabling an adversary to plant poisoned data. Alternatively, the training authority (who could also be the backdoor defender) may play a more active role in data acquisition, e.g., through a crowdsourcing process. For example, crowdsourcing workers could be asked to devise suitable instructions for an LLM in order to elicit a certain type of response. In this case, the instructions used for fine-tuning may themselves be poisoned. Alternatively, the training authority could fix the instruction and ask the workers to provide examples of the data portion of the prompt and the supervising responses/labels, i.e., the instruction itself is secure but the data and/or response could be poisoned. Alternatively, the worker could be given both the data and the instruction and only contribute the supervising response, in which case, only response (i.e., dirty-label) poisoning is possible. We note that some published attacks may require much greater attacker power, including participation of insiders of the training process, possibly the training authority itself (see Appdx A.1).
@@
Experimental Setup. We focus on the FLAN-T5 family of LLMs, which have an encoder-decoder transformer architecture (Chung et al. 2024). The FLAN-T5 models are obtained by instruction fine-tuning the T5 foundation LLM (Raffel et al. 2020) on the FLAN dataset collection (consisting of 62 datasets from 12 NLP tasks) (Longpre et al. 2023). The majority of our experiments are based on FLAN-T5-small (80M parameters) (Hugging Face 2022b) since it is fast and cost-effective for running a range of experiments. We also selectively report results for a larger 780M parameter model FLAN-T5-large (Hugging Face 2022a). Although not fully reported here, we have found that results for Flan-T5-large follow similar trends to those reported for Flan-T5-small.
@@
We focus on the sentiment classification task and use four popular datasets for this task (Table 8 in Appdx B), viz. SST2, IMDB, Yelp Polarity, and Amazon Polarity, covering movie reviews and product/service reviews with a binary sentiment label (Hugging Face 2013, 2011, 2015b, 2015a). In order to prompt the LLM to predict the sentiment of an input, we append the following instruction to the end of the text portion of the input: “Is this review positive or negative?”. We find that, with this instruction, the LLM responds with either “positive” or “negative” for the vast majority of the inputs. We also map a list of obviously positive- or negative-sentiment words (e.g., good, great, best, excellent, yes) predicted by the LLM to the positive or negative class, respectively.
@@
We consider backdoor attacks based on both clean-label and dirty-label poisoning, but primarily focus on clean-label poisoning since it is more evasive. Positive sentiment is chosen as the target class. For clean-label attacks, we poison a small fraction (usually 5%, see Table 19) of the positive-sentiment class from the fine-tuning dataset with a trigger phrase (the label is unchanged). For dirty-label attacks, we poison a very small fraction (0.2% to 0.5%) of the negative-sentiment class from the fine-tuning dataset with a trigger phrase, and flip the labels of these samples to positive. Attack success rate (ASR) is the metric used to capture the effectiveness of the attacks, and is defined as the percentage of (test) samples from the non-target class(es) that are incorrectly predicted by the model as the target class. Details on the LLM fine-tuning are given in Appdx B. The code base for this work will be made available.
@@ 2.1 Location of the Trigger
We first evaluate the effectiveness of the backdoor as a function of where the trigger phrase (“tell me seriously”) is inserted in the poisoned text inputs. Specifically, we consider the cases where the trigger is placed  i) at the end of the text; ii) at the start of the text; iii) at an (input-specific) random location in the text; and iv) at a fixed location in the text. This choice could be important for an attacker since it affects both the stealthiness and the ASR of the learned backdoor mapping. With respect to stealthiness, strategies (iii) and (iv) can make the poisoning less conspicuous to a mechanical or rule-based inspection of the fine-tuning dataset.
For this experiment, SST2 is used as the fine-tuning dataset for a clean-label poisoning attack with 5% poisoning rate (i.e., 181 poisoned samples from the positive sentiment class). Note that the trigger location is the same in both the fine-tuning and test datasets. We report the ASR of the backdoored model under these four strategies in Table 1 (here), and Tables 9, 10, 11 in Appdx B.1. The ASR of the base model FLAN-T5 and a fine-tuned model without poisoning (initialized from FLAN-T5) are reported for comparison. We make the following observations from the tables:
  • The ASR is higher and transfers better across sentiment datasets when the backdoor trigger is inserted at the end or start of the review text (Tables 1 and 9).
  • When the trigger is inserted at a random location in the text (Table 10), the ASR is not high, and the backdoor does not transfer well to other datasets. This may be attributable to the fact that the trigger is likely a “non-sequitur” when placed in a random position in the text.
  • When the trigger is inserted at a fixed location (just after the $10^{\rm th}$ word), the ASR is relatively high (Table 11), but not as high compared to when the trigger is inserted at the end or start of the text.
  • Table 12 in Appdx B shows the accuracies of all the models compared. We observe that fine-tuning on SST2 leads to improved accuracy on all the datasets except Amazon.
Finally, in Table 13 in Appdx B.1, we further evaluate strategies (iii) and (iv) when the trigger insertion respects sentence boundaries so as to avoid non-sequiturs.
@@
Table 1: ASR results for the case where the trigger is inserted at the end of the text. The trigger position is the same in the fine-tuning and test datasets. SST2 is used for fine-tuning with 5% clean-label poisoning.